Eight times a certain number is twenty more than four times the number. Find the number.
Choices:
a. 3
b. 4
c. 5
d. 6

Answer by Cromlix(4381)

Eight times a certain number = 8x
Twenty more than four times a number = 20 + 4x
8x = 20 + 4x
8x - 4x = 20
4x = 20
x = 5
This is the answer (c)
Hope this helps.
:-)